OpenStudy (anonymous):

help please...how to solve e^(8x + 6) − 4 = 0 with an exact answer? do i use ln to solve it?

OpenStudy (anonymous):

ughh deleted again!!!

OpenStudy (anonymous):

\[e^{8x+6}=4\] \[8x+6=\ln(4)\] \[8x=\ln(4)-6\] \[x=\frac{\ln(4)-6}{8}\]

OpenStudy (anonymous):

thanks :D